# Convergence Analysis of Non-Strongly-Monotone Stochastic Quasi-Variational Inequalities

### v2

@@ 1 Introduction
Variational Inequality (VI) problems find applications in various areas like convex Nash games, traffic equilibrium, and economic equilibrium [11]. The Stochastic Variational Inequality (SVI) extends VI theory to address decision-making problems involving uncertainty [13]. An extension of VIs, known as Quasi-Variational Inequality (QVI), emerges when the constraint set depends on the decision variable. From a Nash equilibrium game perspective, QVI captures the interdependencies among players’ strategy sets, particularly in scenarios involving shared resources.
@@
where $K:X\to 2^{X}$ is a set-valued mapping with non-empty, closed and convex values such that $K(x)\subseteq X$ for all $x\in X$, $X\subseteq\mathbb{R}^{n}$ is a convex and compact set, $F(x)\triangleq\mathbb{E}[G(x,\xi)]$, $\xi:\Omega\to\mathbb{R}^{d}$, ${G}:X\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$, and the associated probability space is denoted by $(\Omega,{\cal F},\mathbb{P})$.
@@ 2.1 Generalized Nash Equilibruim
Nash equilibrium (NE) is a fundamental concept in game theory where a finite collection of selfish agents compete with each other and seek to optimize their own individual objectives. An NE is described as a collection of specific strategies chosen by all the players, where no player can reduce their cost by unilaterally changing their strategy within their feasible strategy set. The Generalized Nash equilibrium Problem (GNEP) is an extension of the classic NE Problem (NEP), in which players interact also at the level of the feasible sets, i.e., each player’s strategy set depends on other players’ strategies. This situation arises naturally if the players share some common resource, e.g., a communication link, an electrical transmission line, etc. GNEP has been extensively utilized in the literature to formulate applications arising in economics, operations research, and other fields [9, 20]. Consider $N$ players each with cost function $f_{i}(x_{i},x_{(-i)})\triangleq\mathbb{E}[h(x_{i},x_{(-i)},\xi)]$ for $i=1\ldots,N$, where $x_{i}$ is the strategy of player $i$ and $x_{(-i)}$ is the strategy of other players. Each player $i$’s objective is to solve the following optimization problem:
@@
where $K_{i}(x_{(-i)})=\{x_{i}\in\mathbb{R}^{n_{i}}|g_{i}(x_{i},x_{(-i)})\leq 0\}$ is a closed convex set-valued map. $f_{i},g_{i}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ are continuously differentiable. By defining $K(x)=\prod_{i=1}^{N}K_{i}(x_{(-i)})$ and $F(x)=[\nabla_{x_{i}}f_{i}(x)]_{i=1}^{N}$, finding a GNE will be equivalent to solving the following SQVI problem. $\langle F(x^{*}),x-x^{*}\rangle\geq 0,\forall x\in K(x^{*})$.
@@ 4.2 Assumptions and Technical Lemmas
An operator $F$ has a quadratic growth (QG) property on set $X$ if there exists a constant ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mu_{F}}>0$ such that for any $x\in X$ and $y=\mathcal{P}_{X^{*}}(x)$
@@
It is worth noting that, unlike the strong monotonicity assumption, the QG property does not imply a unique solution. In fact, QG property is a weaker assumption than strong monotonicity [27]. As an example of a QG operator, consider function $f(x)\triangleq g(Ax)+c^{T}x$, where $g$ is a smooth and strongly convex function, $A\in\mathbb{R}^{n\times m}$ is a nonzero general matrix and $c\in\mathbb{R}^{n}$. One can show that $\nabla f(x)$ satisfies the QG property [27] while $\nabla f(x)$ may not be strongly monotone unless $A$ has a full column rank.
@@
(i) The set of optimal solution, $X^{*}$ is nonempty. (ii) Operator $F:X\rightarrow\mathbb{R}^{n}$ is monotone, i.e., $\langle F(x)-F(y),x-y\rangle\geq 0$ for all $x,y\in X$, and satisfies the QG property. (iii) $F$ is $L$-Lipschitz continuous on $X$, i.e.,
@@
If $\mathcal{F}_{k}$ denotes the information history at epoch $k$, then we have the following requirements on the associated filtrations.
@@
Gap Function. Now, we define a gap function to measure the quality of the solution obtained from the algorithm. In particular, for a given iterate $x$ we use $\text{dist}(x,X^{*})\triangleq\|x-\mathcal{P}_{X^{*}}(x)\|$ to find the distance of the solution obtained by the algorithm from the optimal solution set $X^{*}$. Moreover, we call $x$ to be an $\epsilon$-solution if $\|x-\bar{x}\|\leq\epsilon$ where $\bar{x}\triangleq\mathcal{P}_{X^{*}}(x)$.
@@ 5 Proposed Method
There exists $\gamma>0$ such that $\|\mathbf{P}_{K(x)}[u]-\mathbf{P}_{K(y)}[u]\|\leq\gamma\|x-y\|$ for all $x,y,u\in X$.
@@
$\mathbf{P}_{K(x)}[u]=\mathbf{P}_{m(x)+K}[u]=m(x)+\mathbf{P}_{K}[u-m(x)],\quad\forall x,u\in X.$
@@
$\|\mathbf{P}_{K(x)}[u]-\mathbf{P}_{K(y)}[u]\|$ $=\|m(x)-m(y)+\mathbf{P}_{K}[u-m(x)]-\mathbf{P}_{K}[u-m(y)]\|$
$\leq\|m(x)-m(y)\|+\|\mathbf{P}_{K}[u-m(x)]-\mathbf{P}_{K}[u-m(y)]\|$
$\leq 2\|m(x)-m(y)\|\leq 2\tilde{\gamma}.$
This shows that the Assumption 3 holds. Furthermore, from an algorithmic perspective to ensure convergence guarantee due to changes in the constraint set, one approach is to introduce a retraction step [26], denoted as $(1-\alpha)x_{k}+\alpha u_{k}$ for some $\alpha\in[0,1]$. By integrating these ideas, we present the following variant of the SEG method for solving (SQVI).
@@
Input: $x_{0}\in X$, $\eta>0$, $\{N_{k}\}_{k}$, $\{t_{k}\}_{k}$, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{b_{k}\}_{k}}$, $\{\alpha_{k}\}_{k}$ and Algorithm $\mathcal{M}$ satisfying Assumption 4;
@@ 5.1 Convergence Analysis
Next, we introduce a crucial lemma for our convergence analysis. As previously discussed, the problem is not strong monotone and may not possess a unique solution. Therefore, we define the gap function as $\text{dist}(x,X^{*})\triangleq\|x-\bar{x}\|$, where $\bar{x}=\mathcal{P}_{X^{*}}(x)$. Since the optimal solutions are not explicitly available, we need to express $\bar{x}$ based on its first-order optimality condition. This representation will be utilized in the subsequent convergence analysis of the algorithm.
@@
Let $X^{*}$ denote the set of optimal solutions of problem (SQVI). Moreover, for any $x\in X$ define $\bar{x}\triangleq\mathcal{P}_{X^{*}}(x)$. Then, $\bar{x}$ satisfies the following
$\bar{x}=\mathcal{P}_{K(\bar{x})}(\bar{x}-\eta F(\bar{x})),$ (4)
for any $\eta>0$.
@@
At each iteration of $k\geq 0$, we define the error of sample operator $F$ as
@@
For any $k\geq 0$, we define $\bar{x}_{k}\triangleq\mathcal{P}_{X^{*}}(x_{k})$ where $X^{*}$ denotes the set of optimal solutions of problem (SQVI). From Lemma 2 we conclude that $\bar{x}_{k}=\mathcal{P}_{K(\bar{x}_{k})}[\bar{x}_{k}-\eta F(\bar{x}_{k})]$. Using the update rule of $x_{k+1}$ in Algorithm 1 and the fact that $e_{k}$ denotes the error of computing the projection operator, we obtain the following.
@@
Next, let’s examine a scenario in which a constant (mini) batch-size sampling operator of $F$ is accessible during each iteration $k\geq 0$. This configuration arises in various contexts, including online optimization, where problem information becomes progressively available over time. (See Appendix for proof.)
@@
In the following corollary, we demonstrate an improvement in the oracle complexity for the deterministic case, reducing it to $\mathcal{O}(\log(1/\epsilon))$. The proof is included in the appendix and follows a similar approach to that of Corollary 1, with $w$ and $w^{\prime}$ both being zero.
@@ 5.2 Gradient Approach
A natural question that arises after proposing an extra gradient method is whether one can achieve similar convergence results for a gradient method. Therefore, in this section, we shift our focus to gradient method for solving the (SQVI) problem. i.e., letting retraction parameter $b_{k}=0$ in Algorithm 1. In particular, we propose an inexact Gradient SQVI (iG-SQVI) method in Algorithm 2.
@@
Input: $x_{0}\in X$, $\eta>0$, $\{N_{k}\}_{k}$, $\{t_{k}\}_{k}$, $\{\alpha_{k}\}_{k}$ and Algorithm ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}$ satisfying Assumption 4;
@@
$x_{k+1}$;
In the next theorem and corollary, we show the convergence results of Algorithm 2 which is similar to the rate results of Algorithm 1. (See Appendix for proof.)
@@
Let $\{x_{k}\}_{k\geq 0}$ be the iterates generated by Algorithm 2 using step-size $\eta>0$ satisfying $|\eta-\tfrac{\mu_{F}}{L^{2}}|<{\frac{\sqrt{\mu_{F}^{2}-L^{2}(2\gamma-\gamma^{2})}}{L^{2}}}$ and retraction parameter $\alpha_{k}=\bar{\alpha}\in(0,1)$ for $k\geq 0$. SSuppose Assumptions 1-3 hold and $\gamma+\sqrt{1-{\mu_{F}}^{2}/L^{2}}<1$, then for any $T\geq 1$ we have that

### v3

@@ Abstract
keywords: Quasi-Variational Inequalities Stochastic Optimization Generalized Nash Equilibrium Extra-Gradient Method
@@ 1 Introduction
Variational Inequality (VI) problems find applications in various areas like convex Nash games, traffic equilibrium, and economic equilibrium [13]. The Stochastic Variational Inequality (SVI) extends VI theory to address decision-making problems involving uncertainty [45, 16, 24, 19, 1]. An extension of VIs, known as Quasi-Variational Inequality (QVI), emerges when the constraint set depends on the decision variable. From a Nash equilibrium game perspective, QVI captures the interdependencies among players’ strategy sets, particularly in scenarios involving shared resources.
@@
where $K:X\to 2^{X}$ is a lower-hemicontinous set-valued mapping with non-empty, closed and convex values such that $K(x)\subseteq X$ for all $x\in X$, $X\subseteq\mathbb{R}^{n}$ is a convex and compact set, $F(x)\triangleq\mathbb{E}[G(x,\xi)]$, $\xi:\Omega\to\mathbb{R}^{d}$, ${G}:X\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$, and the associated probability space is denoted by $(\Omega,{\cal F},\mathbb{P})$.
@@ 2.1 Generalized Nash Equilibruim
Nash equilibrium (NE) is a fundamental concept in game theory where a finite collection of selfish agents compete with each other and seek to optimize their own individual objectives. An NE is described as a collection of specific strategies chosen by all the players, where no player can reduce their cost by unilaterally changing their strategy within their feasible strategy set. The Generalized Nash equilibrium Problem (GNEP) is an extension of the classic NE Problem (NEP), in which players interact also at the level of the feasible sets, i.e., each player’s strategy set depends on other players’ strategies. This situation arises naturally if the players share some common resource, e.g., a communication link, an electrical transmission line, etc. GNEP has been extensively utilized in the literature to formulate applications arising in economics, operations research, and other fields [11, 25]. Consider $N$ players each with cost function $f_{i}(x_{i},x_{(-i)})\triangleq\mathbb{E}[h(x_{i},x_{(-i)},\xi)]$ for $i=1\ldots,N$, where $x_{i}$ is the strategy of player $i$ , $x_{(-i)}$ is the strategy of other players, and $\xi:\Omega\to\mathbb{R}^{d}$ is a random variable defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$. Each player $i$’s objective is to solve the following optimization problem:
@@
where $K_{i}(x_{(-i)})=\{x_{i}\in\mathbb{R}^{n_{i}}|g_{i}(x_{i},x_{(-i)})\leq 0\}$ is a closed convex set-valued map. $f_{i},g_{i}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ are continuously differentiable. By defining $K(x)=\prod_{i=1}^{N}K_{i}(x_{(-i)})$ and $F(x)=[\nabla_{x_{i}}f_{i}(x)]_{i=1}^{N}$, finding a GNE will be equivalent to solve the following SQVI problem: $\langle F(x^{*}),x-x^{*}\rangle\geq 0,\forall x\in K(x^{*})$.
@@ 4.2 Assumptions and Technical Lemmas
An operator $F$ has a quadratic growth (QG) property on set $X$ if there exists a constant ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mu_{F}}>0$ such that for any $x\in X$ and some $y\in\mathcal{P}_{X^{*}}[x]$
@@
It is worth noting that, unlike the strong monotonicity assumption, the QG property does not imply a unique solution. In fact, QG property is a weaker assumption than strong monotonicity [6, 32]. As an example of a QG operator, consider function $f(x)\triangleq g(Ax)+c^{T}x$, where $g$ is a smooth and strongly convex function, $A\in\mathbb{R}^{n\times m}$ is a nonzero general matrix and $c\in\mathbb{R}^{n}$. One can show that $\nabla f(x)$ satisfies the QG property [32] while $\nabla f(x)$ may not be strongly monotone unless $A$ has a full column rank. See [6] and [32] for more examples.
@@
(i) The set of optimal solution $X^{*}$ is nonempty. (ii) Operator $F:X\rightarrow\mathbb{R}^{n}$ is monotone, i.e., $\langle F(x)-F(y),x-y\rangle\geq 0$ for all $x,y\in X$, and satisfies the QG property. (iii) $F$ is $L$-Lipschitz continuous on $X$, i.e.,
@@
If $\mathcal{F}_{k}$ denotes the information history at epoch $k$, then we have the following requirements on the associated filtration which imposes unbiasedness and bounded variance conditions on the stochastic error in estimating $F(x_{k})$ using a finite sample average.
@@
Next, we define a gap function to measure the quality of the solution obtained from the algorithm.
Definition 2 (Gap function.).
For a given iterate $x$ we use $\text{dist}(x,X^{*})\triangleq\|x-\mathcal{P}_{X^{*}}(x)\|$ to find the distance of the solution obtained by the algorithm from the optimal solution set $X^{*}$. Moreover, we call $x$ to be an $\epsilon$-solution if $\|x-\bar{x}\|\leq\epsilon$ where $\bar{x}\in\mathcal{P}_{X^{*}}(x)$.
@@ 5 Proposed Method
There exists $\gamma>0$ such that $\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K(x)}[u]-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K(y)}[u]\|\leq\gamma\|x-y\|$ for all $x,y,u\in X$.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K(x)}[u]={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{m(x)+K}[u]=m(x)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K}[u-m(x)],\quad\forall x,u\in X.$
@@
$\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K(x)}[u]-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K(y)}[u]\|$ $=\|m(x)-m(y)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K}[u-m(x)]-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K}[u-m(y)]\|$
$\leq\|m(x)-m(y)\|+\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K}[u-m(x)]-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K}[u-m(y)]\|$
$\leq 2\|m(x)-m(y)\|\leq 2\tilde{\gamma}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|x-y\|}.$
This shows that Assumption 3 holds. Another important class of problems satisfying Assumption 3 is when the solution set $K(x)$ is a bounded polyhedral moving set with right-hand-side linearly dependent on $x$. In such cases, by Corollary 2 in [5], the projection map $\mathcal{P}_{K(x)}$ is Lipschitz continuous. Furthermore, from an algorithmic perspective to ensure convergence guarantee due to changes in the constraint set, one approach is to introduce a retraction step [31], denoted as $(1-\alpha)x_{k}+\alpha u_{k}$ for some $\alpha\in[0,1]$. By integrating these ideas, we present the following variant of the SEG method for solving (SQVI).
@@
Input: $x_{0}\in X$, $\eta>0$, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T>0}$, $\{N_{k}\}_{k}$, $\{t_{k}\}_{k}$, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{b_{k}\}_{k}}$, $\{\alpha_{k}\}_{k}$ and Algorithm $\mathcal{M}$ satisfying Assumption 4;
@@ 5.1 Convergence Analysis
Next, we introduce a crucial lemma for our convergence analysis. As previously discussed, the problem is not strong monotone and may not possess a unique solution (see [12, 14]). Therefore, we define the gap function as $\text{dist}(x,X^{*})\triangleq\|x-\bar{x}\|$, where $\bar{x}\in\mathcal{P}_{X^{*}}(x)$. Since the optimal solutions are not explicitly available, we need to express $\bar{x}$ based on its first-order optimality condition. This representation will be utilized in the subsequent convergence analysis of the algorithm.
@@
Let $X^{*}$ denote the set of solutions of problem (SQVI). Then, for any $\bar{x}\in X^{*}$ and $\eta>0$ the following holds
$\bar{x}=\mathcal{P}_{K(\bar{x})}(\bar{x}-\eta F(\bar{x})).$ (4)
@@
At each iteration $k\geq 0$, we define the error of sample operator $F$ as
@@
For any $k\geq 0$, we define $\bar{x}_{k}\in\mathcal{P}_{X^{*}}(x_{k})$ where $X^{*}$ denotes the set of solutions of problem (SQVI). From Lemma 2 we conclude that $\bar{x}_{k}=\mathcal{P}_{K(\bar{x}_{k})}[\bar{x}_{k}-\eta F(\bar{x}_{k})]$. Using the update rule of $x_{k+1}$ in Algorithm 1 and the fact that $e_{k}$ denotes the error of computing the projection operator, we obtain the following.
@@
Next, let’s examine a scenario in which a constant (mini) batch-size sampling operator of $F$ is accessible during each iteration $k\geq 0$. This configuration arises in various contexts, including online optimization, where problem information becomes progressively available over time. (See Appendix Section A for the proof.)
@@
In the following corollary, we demonstrate an improvement in the oracle complexity for the deterministic case, reducing it to $\mathcal{O}(\log(1/\epsilon))$. The proof is provided in Appendix Section A and follows a similar approach to that of Corollary 1, with $w$ and $w^{\prime}$ both being zero.
@@ 5.2 Gradient Approach
A natural question that arises after proposing an extra gradient method is whether one can achieve similar convergence results for a gradient method. Therefore, in this section, we shift our focus to gradient method for solving the (SQVI) problem, i.e. letting retraction parameter $b_{k}=0$ in Algorithm 1. In particular, we propose an inexact Gradient SQVI (iG-SQVI) method in Algorithm 2.
@@
Input: $x_{0}\in X$, $\eta>0$, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T>0}$, $\{N_{k}\}_{k}$, $\{t_{k}\}_{k}$, $\{\alpha_{k}\}_{k}$ and Algorithm ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}$ satisfying Assumption 4;
@@
$x_{k+1}$.
In the next theorem and corollary, we show the convergence results of Algorithm 2 which is similar to the rate results of Algorithm 1. (See Appendix Section B for the proof.)
@@
Let $\{x_{k}\}_{k\geq 0}$ be the iterates generated by Algorithm 2 using step-size $\eta>0$ satisfying $|\eta-\tfrac{\mu_{F}}{L^{2}}|<{\frac{\sqrt{\mu_{F}^{2}-L^{2}(2\gamma-\gamma^{2})}}{L^{2}}}$ and retraction parameter $\alpha_{k}=\bar{\alpha}\in(0,1)$ for $k\geq 0$. Suppose Assumptions 1-3 hold and $\gamma+\sqrt{1-{\mu_{F}}^{2}/L^{2}}<1$, then for any $T\geq 1$ we have that
@@
Note that to obtain convergence rate guarantees, we required the condition $\gamma+\sqrt{1-\frac{\mu_{F}^{2}}{L^{2}}}<1$, which is problem-dependent and commonly appears in prior works on deterministic QVI problems with convergence guarantees (e.g., see Theorem 2 and 3 in [31] and Theorem 5 in [35]). Whether convergence rate guarantees—such as sublinear or linear rates—can be established without relying on this condition remains an open question, particularly due to the added complexity introduced by feasible sets that depend on the decision variable. In the next section, we show that almost sure convergence can be achieved without relying on this structural condition.
5.3 Almost Sure Convergance
To establish almost sure convergence of the iterates, we first state the following Robbins-Siegmund lemma (see Lemma 11 in [41]).
Lemma 3 (Robbins-Siegmund).
Suppose $\omega_{t},\eta_{t},\upsilon_{t}$ and $\psi_{t}$ are nonnegative random variables that satisfy:
$\mathbb{E}\left[\omega_{t+1}\mid\mathcal{F}_{t}\right]\leq\left(1+\eta_{t}\right)\omega_{t}+\upsilon_{t}-\psi_{t},\quad\forall t\geq 0.$
[…]
Now we are ready to show that the iterates generated by Algorithm 2 have at least one limit point, and every limit point is a solution of the problem SQVI almost surely. (See Appendix Section C for the proof.)
Theorem 3 (Almost sure convergence).
Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by Algorithm 2, and suppose Assumption 1 and 2 hold and $\{N_{k}\}_{k}$ is an increasing sequence, such that $\sum_{k=0}^{\infty}\tfrac{1}{\sqrt{N_{k}}}<\infty$. Then the sequence $\{x_{k}\}$ has at least one limit point, and every limit point is a solution of the SQVI almost surely.